OpenStudy (anonymous):

Simplify 3√5 - 2√7 + √45 -√28. 2√12 2√2 6√5 - 4√7 6√10 - 4√14

OpenStudy (unklerhaukus):

what are prime factors of 45?

OpenStudy (anonymous):

5 , 15, 3

OpenStudy (unklerhaukus):

15 is not prime

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

ok i gt ya

OpenStudy (unklerhaukus):

45 = 3 * 3 * 5 right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

you have one term that is √45 √45 = √(3 * 3 * 5) = √(3^2 * 5) = √3^2 * √5 = 3 * √5

OpenStudy (unklerhaukus):

because √x^2 = x and √(x * y) = √x * √y

OpenStudy (unklerhaukus):

do you know the prime factorization of 28?

OpenStudy (anonymous):

\[3\sqrt{5}-2\sqrt{7}+3\sqrt{5}+\sqrt{9}\sqrt{5}-2\sqrt{7}\]\[3\sqrt{5}+3\sqrt{5}-2\sqrt{7}-2\sqrt{7}\]\[6\sqrt{5}-4\sqrt{7}\]

OpenStudy (anonymous):